OpenStudy (anonymous):

If a ball is thrown directly upward with a velocity of 46 ft/s, its height (in feet) after t seconds is given by y = 46t − 16t2. What is the maximum height attained by the ball? (Round your answer to the nearest whole number.)

OpenStudy (confluxepic):

\[\huge \color{blue}{Welcome~To~OpenStudy!!!}\]

OpenStudy (anonymous):

You have to find the vertex (highest point in the graph)

OpenStudy (anonymous):

the x-coordinate of the vertex is -b ---- 2a in the equation ax^2+bx+c=0 where a=-16 b=46 c-0

OpenStudy (anonymous):

After you find that x-value, then you plug it into -16t^2+46 and that will give you the max height

OpenStudy (anonymous):

for the \[\frac{ -b }{ 2a }\] i got \[\frac{ 23 }{ 16 }\]

OpenStudy (anonymous):

Alright, now you have to plug that into -16t^2+46t

OpenStudy (anonymous):

you have to plug in 23/16 for everywhere you see t

OpenStudy (anonymous):

yea, It should be 33.0625

OpenStudy (anonymous):

I must have done something wrong because my end answaer was 529.

OpenStudy (anonymous):

How did you get to 33.0625 from 529